OpenStudy (anonymous):

Evaluate b^2c^-1 for b = 8 and c = -4. i got 256.. is that right?

OpenStudy (anonymous):

Check once again..

OpenStudy (campbell_st):

this is \[(8)^2\times \frac{1}{-4} = 64 \times -\frac{1}{4}\]

OpenStudy (anonymous):

@waterineyes is it 1/16?

OpenStudy (campbell_st):

\[8 \times 8 \times -\frac{1}{4}\]

OpenStudy (anonymous):

?

OpenStudy (campbell_st):

well what if I write it this way \[-\frac{ 8 \times 8}{4} =\]

OpenStudy (anonymous):

-16?

OpenStudy (campbell_st):

correct

OpenStudy (anonymous):

so is that the final answer?

OpenStudy (anonymous):

thanks

OpenStudy (campbell_st):

yes,,, provided you question was correctly written